OpenStudy (anonymous):

Find the 32nd term of the sequence: 9, 4, -1, -6....

hero (hero):

Just keep adding -5

OpenStudy (anonymous):

so would it be -200 ?

OpenStudy (anonymous):

a = 9, d = -5 32nd term = 9 +31(-5) =

OpenStudy (anonymous):

how did you get -200?

OpenStudy (anonymous):

Educate_me, there is a difference of -5 between every term and the next. There are 31 such steps between the first term and the 32nd. Therefore the 32nd term is 9 - 31(5) = -146

OpenStudy (anonymous):

im confused.

OpenStudy (anonymous):

what don't you understand? Let's try an example: The 4th term of the sequence = a + (n-1)d = a + (4-1)d = a + 3d = 9 + 3(-5) = 9 - 15 = -6

OpenStudy (anonymous):

I'm getting a different answer. You are trying to find the nth term or the 32nd term. So the difference is always -5. So your thingy will be -5n + (some number). Remember that number doesn't always work in the first and second case as is here.

OpenStudy (anonymous):

After I did it again I got -146 ,

OpenStudy (anonymous):

Yeah, Luminaire is right, I meant I got a different answer from Educateme. My method is -5n+14

OpenStudy (anonymous):

Thaks guys... but one more question.. Write a recursive and an explicit formula for each sequence. 0, 6, 12, 18, 24 and 27, 15, 3, -9, -21

OpenStudy (anonymous):

Do as I said earlier (remember in some cases it doesn't work in the first one or two cases). In the first one the difference is always 6. Therefore 6n + (some number). In each case you test, for the first case is 6 (1) + some number. What number added to 6 gives the next number and on.

OpenStudy (anonymous):

but , how could we form a formula. the first sequence the common difference is 6and the first term is 0 ? so woud the formula be an=a1 (n-1) (6) ?

OpenStudy (anonymous):

It doesn't say if you starting from 0 or 1. Let's assume you're starting at 1. It is 6n + 6. Or 6(1) + 6 which is 12, but remember, it doesn't always work in the first one or two terms. Just continue like that.

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

Actually, it should have been obvious to me the first one you gave starts at the zeroth term and works perfectly throughout 6n+6. See if you can do the second one.